OpenStudy (anonymous):

Convert the radian measure to degree measure. Use the value of π found on a calculator and round answers to two decimal places. negative nineteen divided by six, pi

OpenStudy (anonymous):

\[-\frac{ 19 }{ 6 } \pi \times \frac{ 180 }{ \pi }\]

OpenStudy (anonymous):

thanks!